# Estimating invertible processes in Hilbert spaces, with applications to functional ARMA processes

### v2

@@ 1 Introduction
The fMA and fARMA cases have received less attention. Turbillon et al. (2008) estimated the fMA$(1)$ operator assuming commutativity, and Kuenzer (2024) derived consistent estimators for fARMA operators using a two-step approach, involving sequential AR and MA estimation with dimension reduction via principal components. Central limit theorems for fAR operator estimates are derived in Mas and Pumo (2009).
This article formulates consistency results with quantifiable rates for infinite-dimensional operators defining general $\mathcal{H}$-valued invertible processes with white noise innovations, reyling on Yule–Walker equations and a Tikhonov-regularized estimator, avoiding assumptions tied to specific models. The consistency results are established under a Sobolev-type condition and assuming general consistency rates for certain lagged covariance operator estimates that hold in many standard situations, such as under $L^{p}$-$m$-approximability, a weak dependence concept for fTSA introduced in Hörmann and Kokoszka (2010). We apply these findings to derive consistency for the infinite-dimensional operators defining causal linear processes, including fAR, fMA, and fARMA processes with arbitrary AR and MA orders. Notably, our results are derived under mild conditions, permitting multi-dimensional eigenspaces of the covariance operator, and without requiring specific rates of eigenvalue decay. Provided the operators exhibit a sufficiently high degree of smoothness and the eigenvalue decay of the error covariance operators is slow, our approach yields the near-parametric rate $N^{-1/2}.$
@@ 2.1 Notation
The additive identity in linear spaces is denoted by 00, and the identity map by $\mathbb{I}$. In Cartesian product spaces $V^{n}$, $n\in\mathbb{N}$, scalar multiplication and addition are defined component-wise. Let $(\mathcal{B},\|\cdot\|)$ be a Banach space and $(\mathcal{H},\langle\cdot,\cdot\rangle)$, $(\mathcal{H}_{\star},\langle\cdot,\cdot\rangle_{\star})$ Hilbert spaces with their respective norms and inner products. Then, $\mathcal{B}^{n}$ with norm $\|x\|^{2}=\sum_{i=1}^{n}\|x_{i}\|^{2}$ is a Banach space, and $\mathcal{H}^{n}$ with inner product $\langle x,y\rangle=\sum_{i=1}^{n}\langle x_{i},y_{i}\rangle$ is a Hilbert space. The spaces of bounded linear, Hilbert–Schmidt (H-S), and nuclear (trace-class) operators $A\colon\mathcal{H}\to\mathcal{H}_{\star}$ are denoted by $\mathcal{L}_{\mathcal{H},\mathcal{H}_{\star}}$, $\mathcal{S}_{\mathcal{H},\mathcal{H}_{\star}}$, and $\mathcal{N}_{\mathcal{H},\mathcal{H}_{\star}}$, respectively, with norms $\|\cdot\|_{\mathcal{L}}$, $\|\cdot\|_{\mathcal{S}}$, and $\|\cdot\|_{\mathcal{N}}$, and H-S inner product $\langle\cdot,\cdot\rangle_{\mathcal{S}}$. For $\mathcal{T}\in\{\mathcal{L,S,N}\}$, set $\mathcal{T_{H}}\coloneqq\mathcal{T}_{\mathcal{H,H}}$. The projection onto the closed subspace spanned by $a_{m},\dots,a_{n}$ of a complete orthonormal system (CONS) $(a_{j})_{j\in\mathbb{N}}$ in $\mathcal{H}$ is denoted $\coprod^{a_{n}}_{a_{m}}\colon\mathcal{H}\to\mathrm{span}\{a_{m},\dots,a_{n}\}\subset\mathcal{H}$. All random variables are defined on a common probability space $(\Omega,\mathfrak{A},\mathbb{P})$. For a sequence $(c_{n})_{n\in\mathbb{N}}\subset(0,\infty)$, $X_{n}=\mathrm{o}_{\mathbb{P}}(c_{n})$ and $X_{n}=\mathrm{O}_{\mathbb{P}}(c_{n})$ as $n\to\infty$ denote $X_{n}/c_{n}$ converges in probability and is stochastically bounded, respectively. The space of random variables $X\in\mathcal{H}$ with $\mathbb{E}\|X\|^{p}<\infty,$ $p\in[1,\infty)$, is denoted by $L^{p}_{\mathcal{H}}=L^{p}_{\mathcal{H}}(\Omega,\mathfrak{A},\mathbb{P})$. A sequence $(X_{k})\subset L^{1}_{\mathcal{H}}$ is centered if $\mathbb{E}(X_{k})=0$ for all $k$ with expectation understood as a Bochner integral. For sequences $(a_{n})_{n\in\mathbb{N}},(b_{n})_{n\in\mathbb{N}}\subset\mathbb{R}$, $a_{n}\asymp b_{n}$ as $n\to\infty$ indicates asymptotic equivalence up to constant factors. Finally, $\mathds{1}_{A}(x)$ denotes the indicator function of a set $A,$ and $\delta_{ij}$ the Kronecker delta. Unless stated otherwise, asymptotic statements refer to $N\to\infty$.
@@ 3 Estimation of the operators in invertible processes
$(X_{k})_{k\in\mathbb{Z}}\subset L^{2}_{\mathcal{H}}$ is a centered, invertible process with representation (3.1), and the operators $\psi_{j}\in\mathcal{S}_{\mathcal{H}}$.
@@
where ϕsubscriptitalic-$\phi_{\ell}\in\mathcal{L_{H}}$ with $\sum^{\infty}_{\ell=-\infty}\|\phi_{\ell}\|_{\mathcal{L}}<\infty.$ Then, Assumption 3.3 is satisfied if the linear process operators fulfill the stronger condition $\sum^{\infty}_{k=1}\sum^{\infty}_{\ell=k}\left[\|\phi_{\ell}\|_{\mathcal{L}}+\|\phi_{-\ell}\|_{\mathcal{L}}\right]<\infty.$
@@
$S_{\Psi_{\!L}}\!(\beta)\coloneqq\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle\Psi_{\!L},c_{i}\otimes d_{j}\rangle^{2}_{\mathcal{S}}\,(1+i^{2\beta}\!+j^{2\beta})<\infty.$
@@
For all $N$ sufficiently large, with probability one the eigenvalues $\hat{\lambda}_{j}=\hat{\lambda}_{j}(L)$ of $\hat{\mathscr{C}}_{\!\boldsymbol{X}^{[L]}}$ satisfy $\hat{\lambda}_{j}\neq\hat{\lambda}_{j+1}$ for all $j\in\{1,...,K\}$ and $L.$
@@ 5 Consequences for fARMA processes
The asymptotic results in Section 3 extend to the estimation of fAR, fMA, and fARMA parameters. These models are extensively studied in Bosq (2000), particularly the fAR case, while Spangenberg (2013) provides general conditions for causal and invertible representations in Banach spaces. Estimation results for the fAR(1) operator are available in Bosq (2000) and Caponera and Panaretos (2022). Additionally, Turbillon et al. (2008) estimated the fMA(1) operator assuming commutativity with the error covariance operator, and Aue and Klepsch (2017); Klepsch et al. (2017) studied prediction for fARMA$(p,q)$ processes. Kuenzer (2024) derived consistent fMA(1) and fARMA(1,1) operator estimates using an adaptation of the Hannan–Rissanen method (Hannan and Rissanen, 1982). By using operator estimates from both the linear and inverted representations, we derive consistent fAR, fMA, and fARMA operator estimates with explicit rates.
@@
A centered process $(X_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is a functional autoregressive process of order $p\in\mathbb{N}$ (fAR$(p)$) if
@@
A centered process $(X_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is a functional moving-average process of order $q\in\mathbb{N}$ (fMA$(q)$) if
@@
We finally turn to the estimation of fARMA processes. A centered process $(X_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is called a functional autoregressive moving-average process of order $p,q\in\mathbb{N}$ (fARMA$(p,q)$) if
@@
where $(\varepsilon_{k})_{k\in\mathbb{Z}}\subset\mathcal{H}$ is a WN, and $\alpha_{i},\beta_{j}\in\mathcal{L_{H}}$ are operators with $\alpha_{p}\neq 0,\beta_{q}\neq 0.$ We combine this fARMA$(p,q)$ representation with its inverted form (3.1). As with multivariate ARMA processes, the operators $\alpha_{i}$ and $\beta_{j}$ are generally not identifiable. Sufficient conditions for identifiability are provided in Kuenzer (2024) (equation (10) and Propositions 1–3). We assume that the fARMA models considered are identifiable and admit invertible, centered, linear process solutions.
To describe the main idea, we first discuss estimation for fARMA(1,1)11(1,1)( 1 , 1 ) processes of the form
@@
Although $\hat{\Psi}_{\!L}$ in (3.10) provides consistent asymptotic estimates for all $\psi_{i}$, consistent estimates for the fARMA(1,1)11(1,1)( 1 , 1 ) operators cannot be directly obtained, as $\psi_{1}\in\mathcal{S_{H}}$ does not possess a bounded inverse. However, using $\hat{\psi}_{i}=\hat{\Psi}^{(i)}_{\!L}$, equation (5.3) leads to the following estimators for $\alpha_{1}$ and $\beta_{1}$:
@@
$\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle\beta_{1},f_{i}\otimes f_{j}\rangle^{2}_{\mathcal{S}}(1+i^{2\gamma}\!+j^{2\gamma})<\infty.$
@@
The dimensions of all eigenspaces of $\psi_{1}\psi^{\ast}_{1}$ are bounded above by some $d\in\mathbb{N}.$
Assumption 5.4.
For all $N$ sufficiently large, with probability one the eigenvalues of $\hat{\psi}_{1}\hat{\psi}^{\ast}_{1}$ satisfy $\hat{\rho}_{j}\neq\hat{\rho}_{j+1}$ for all $j\in\{1,...,M\}.$
@@
Suppose $(X_{k})$ is the fARMA(1,1)11(1,1)( 1 , 1 ) process in (5.2). Further, let the conditions of Theorem 3.1 and Assumptions 5.15.4 hold, and let $M=M_{\!N}\rightarrow\infty$ with $M^{\gamma+1/2}P^{2}_{\!M}=\mathrm{O}(K^{\beta})$ and $\gamma_{N}=\mathrm{O}(\rho_{M}M^{-\gamma}),$ with $P_{M}\coloneqq(\rho_{M}-\rho_{M_{b}})^{-1}$ as in (3.11). Then,
@@
where ^(^^$\boldsymbol{\hat{\prod}}^{\dagger}\coloneqq(\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}+\gamma_{N}\hskip 0.50003pt\mathbb{I})^{-1},$ $(M_{N})\subset\mathbb{N}$ and $(\gamma_{N})\subset(0,\infty)$ are sequences with $M_{N}\to\infty$ and $\gamma_{N}\rightarrow 0,$ and $\hat{h}_{1},\dots,\hat{h}_{\!M}$ are the eigenfunctions associated to the eigenvalues $\hat{\zeta}_{1}\geq\cdots\geq\hat{\zeta}_{M}\geq 0$ of $\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}.$ Further, $\hat{B}_{[i]},$ $B_{[i]},$ $\hat{\Psi}^{\prime\prime\prime}_{[i]},$ $\Psi^{\prime\prime\prime}_{[i]}\in\mathcal{S}_{\mathcal{H}^{\max(i,q)},\mathcal{H}}$ are for each $i,q$ defined by
@@
$\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle B_{q},h_{i}\otimes d_{j}\rangle^{2}_{\mathcal{S}}\,(1+i^{2\gamma}\!+j^{2\gamma})<\infty,$
@@
The dimensions of all eigenspaces of $\boldsymbol{\prod}\boldsymbol{\prod}^{\ast}$ are bounded above by some $d\in\mathbb{N}.$
Assumption 5.8.
For all $N$ sufficiently large, with probability one the eigenvalues of $\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}$ satisfy $\hat{\zeta}_{j}\neq\hat{\zeta}_{j+1}$ for all $j\in\{1,...,M\}.$
@@
Suppose $(X_{k})$ is the fARMA$(p,q)$ process in (5.1). Further, let the conditions of Theorem 3.1 and Assumptions 5.55.8 hold, and let $M=M_{\!N}\to\infty$ with $M^{\gamma+1/2}Z^{2}_{\!M}=\mathrm{O}(K^{\beta})$ and $\gamma_{N}=\mathrm{O}(\zeta_{M}M^{-\gamma})$, where $Z_{\!M}\coloneqq(\zeta_{M}-\zeta_{M_{b}})^{-1}$ as in (3.11), with eigenvalues $\zeta_{j}$ of $\boldsymbol{\prod}\boldsymbol{\prod}^{\ast}$. Then,
@@ 6 Concluding remarks
where $(\varepsilon_{k})$ is i.i.d., $N$ is the sample size, and $k_{N}\to\infty$. This condition is weaker than standard invertibility (3.1), requiring only convergence in probability without assuming finite second moments or constraining $(k_{N})$. Standard invertibility implies mild Granger–Andersen invertibility, but not vice versa. As noted by Granger and Andersen (1978), certain functional time series, such as bilinear processes in separable Hilbert spaces (Hörmann and Kokoszka, 2010), satisfy mild but not standard invertibility. Others, like the non-linear MA$(1)$ process $X_{k}=\alpha_{1}(\varepsilon^{2}_{k-1})+\varepsilon_{k}$ in $L^{2}[0,1]$, may be stationary yet fail to be invertible in both senses.
@@
Acknowledgements
The authors are grateful to the two referees and the associate editor for their constructive comments on an earlier version of this article, which greatly improved the quality of the paper. Parts of the paper were written while Sebastian Kühnert was employed at the University of California, Davis.
Funding
Sebastian Kühnert was partially supported by TRR 391 Spatio-temporal Statistics for the Transition of Energy and Transport (Project number 520388526) funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation). Gregory Rice was partially supported by the Discovery Grant RGPIN 50503-11525 3100 105 from the Natural Science and Engineering Research Council of Canada.
Appendix A Proofs

### v3

@@ 1 Introduction
The fMA and fARMA cases have received less attention. Turbillon et al. (2008) estimated the fMA(1)(1)( 1 ) operator assuming commutativity, and Kuenzer (2024) derived consistent estimators for fARMA operators using a two-step approach, involving sequential AR and MA estimation with dimension reduction via principal components. Central limit theorems for fAR operator estimates are derived in Mas and Pumo (2009).
This article formulates consistency results with quantifiable rates for infinite-dimensional operators defining general $\mathcal{H}$-valued invertible processes with white noise innovations, reyling on Yule–Walker equations and a Tikhonov-regularized estimator, avoiding assumptions tied to specific models. The consistency results are established under a Sobolev-type condition and assuming general consistency rates for certain lagged covariance operator estimates that hold in many standard situations, such as under $L^{p}$-$m$-approximability, a weak dependence concept for fTSA introduced in Hörmann and Kokoszka (2010). We apply these findings to derive consistency for the infinite-dimensional operators defining causal linear processes, including fAR, fMA, and fARMA processes with arbitrary AR and MA orders. Notably, our results are derived under mild conditions, permitting multi-dimensional eigenspaces of the covariance operator, and without requiring specific rates of eigenvalue decay. Provided the operators exhibit a sufficiently high degree of smoothness and the eigenvalue decay of the error covariance operators is slow, our approach nearly achieves the parametric rate $N^{-1/2}.$
@@ 2.1 Notation
The additive identity in linear spaces is denoted by 0, and the identity map by $\mathbb{I}$. In Cartesian product spaces $V^{n}$, $n\in\mathbb{N}$, scalar multiplication and addition are defined component-wise. Let $(\mathcal{B},\|\cdot\|)$ be a Banach space and $(\mathcal{H},\langle\cdot,\cdot\rangle)$, $(\mathcal{H}_{\star},\langle\cdot,\cdot\rangle_{\star})$ Hilbert spaces with their respective norms and inner products. Then, $\mathcal{B}^{n}$ with norm $\|x\|^{2}=\sum_{i=1}^{n}\|x_{i}\|^{2}$ is a Banach space, and $\mathcal{H}^{n}$ with inner product $\langle x,y\rangle=\sum_{i=1}^{n}\langle x_{i},y_{i}\rangle$ is a Hilbert space. The spaces of bounded linear, Hilbert–Schmidt (H-S), and nuclear (trace-class) operators $A\colon\mathcal{H}\to\mathcal{H}_{\star}$ are denoted by $\mathcal{L}_{\mathcal{H},\mathcal{H}_{\star}}$, $\mathcal{S}_{\mathcal{H},\mathcal{H}_{\star}}$, and $\mathcal{N}_{\mathcal{H},\mathcal{H}_{\star}}$, respectively, with norms $\|\cdot\|_{\mathcal{L}}$, $\|\cdot\|_{\mathcal{S}}$, and $\|\cdot\|_{\mathcal{N}}$, and H-S inner product $\langle\cdot,\cdot\rangle_{\mathcal{S}}$. For $\mathcal{T}\in\{\mathcal{L,S,N}\}$, set $\mathcal{T_{H}}\coloneqq\mathcal{T}_{\mathcal{H,H}}$. The projection onto the closed subspace spanned by $a_{m},\dots,a_{n}$ of a complete orthonormal system (CONS) $(a_{j})_{j\in\mathbb{N}}$ in $\mathcal{H}$ is denoted $\coprod^{a_{n}}_{a_{m}}\colon\mathcal{H}\to\mathrm{span}\{a_{m},\dots,a_{n}\}\subset\mathcal{H}$. All random variables are defined on a common probability space $(\Omega,\mathfrak{A},\mathbb{P})$. For a sequence $(c_{n})_{n\in\mathbb{N}}\subset(0,\infty)$, $X_{n}=\mathrm{o}_{\mathbb{P}}(c_{n})$ and $X_{n}=\mathrm{O}_{\mathbb{P}}(c_{n})$ as $n\to\infty$ denote $X_{n}/c_{n}$ converges in probability and is stochastically bounded, respectively. The space of random variables $X\in\mathcal{H}$ with $\mathbb{E}\|X\|^{p}<\infty,$ $p\in[1,\infty)$, is denoted by $L^{p}_{\mathcal{H}}=L^{p}_{\mathcal{H}}(\Omega,\mathfrak{A},\mathbb{P})$. A sequence $(X_{k})\subset L^{1}_{\mathcal{H}}$ is centered if $\mathbb{E}(X_{k})=0$ for all $k$ with expectation understood as a Bochner integral. For sequences $(a_{n})_{n\in\mathbb{N}},(b_{n})_{n\in\mathbb{N}}\subset\mathbb{R}$, $a_{n}\asymp b_{n}$ as $n\to\infty$ indicates asymptotic equivalence up to constant factors. Finally, $\mathds{1}_{A}(x)$ denotes the indicator function of a set $A,$ and $\delta_{ij}$ the Kronecker delta. Unless stated otherwise, asymptotic statements refer to $N\to\infty$.
@@ 3 Estimation of the operators in invertible processes
$(X_{k})_{k\in\mathbb{Z}}\subset L^{2}_{\mathcal{H}}$ is a centered, invertible process with representation (3.1), with operators $\psi_{j}\in\mathcal{S}_{\mathcal{H}}$.
@@
where $\phi_{\ell}\in\mathcal{L_{H}}$ with $\sum^{\infty}_{\ell=-\infty}\|\phi_{\ell}\|_{\mathcal{L}}<\infty.$ Then, Assumption 3.3 is satisfied if the linear process operators fulfill the stronger condition $\sum^{\infty}_{k=-\infty}|k|\,\|\phi_{k}\|_{\mathcal{L}}<\infty.$
@@
$S_{\Psi_{\!L}}\!(\beta)\coloneqq\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle\Psi_{\!L},c_{i}\otimes d_{j}\rangle^{2}_{\mathcal{S}}\,(1+i^{2\beta}\!+\hskip-0.50003ptj^{2\beta})<\infty.$
@@
For all $N$ sufficiently large, with probability one the eigenvalues $\hat{\lambda}_{j}=\hat{\lambda}_{j}(L)$ of $\hat{\mathscr{C}}_{\!\boldsymbol{X}^{[L]}}$ satisfy $\hat{\lambda}_{j}\neq\hat{\lambda}_{j+1}$ for all $j\in\{1,...,K+dL\}$ and $L,$ with $d$ from Assumption 3.5.
@@ 5 Consequences for fARMA processes
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
To describe the main idea, we first discuss estimation for fARMA(1,1)(1,1)( 1 , 1 ) processes of the form
@@
Although $\hat{\Psi}_{\!L}$ in (3.10) provides consistent asymptotic estimates for all $\psi_{i}$, consistent estimates for the fARMA(1,1)(1,1)( 1 , 1 ) operators cannot be directly obtained, as $\psi_{1}\in\mathcal{S_{H}}$ does not possess a bounded inverse. However, using $\hat{\psi}_{i}=\hat{\Psi}^{(i)}_{\!L}$, equation (5.3) leads to the following estimators for $\alpha_{1}$ and $\beta_{1}$:
@@
$\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle\beta_{1},f_{i}\otimes f_{j}\rangle^{2}_{\mathcal{S}}(1+i^{2\gamma}\!+\hskip-0.50003ptj^{2\gamma})<\infty.$
@@
There exists $d\in\mathbb{N}$ such that:
  • (a)
    The dimensions of all eigenspaces of $\psi_{1}\psi^{\ast}_{1}$ are bounded above by $d;$
  • (b)
    For all $N$ sufficiently large, with probability one the eigenvalues of $\hat{\psi}_{1}\hat{\psi}^{\ast}_{1}$ satisfy $\hat{\rho}_{j}\neq\hat{\rho}_{j+1}$ for all $j\in\{1,...,M+d\}.$
@@
Suppose $(X_{k})$ is the fARMA(1,1)(1,1)( 1 , 1 ) process in (5.2). Further, let the conditions of Theorem 3.1 and Assumptions 5.15.3 hold, and let $M=M_{\!N}\rightarrow\infty$ with $M^{\gamma+1/2}P^{2}_{\!M}=\mathrm{O}(K^{\beta})$ and $\gamma_{N}=\mathrm{O}(\rho_{M}M^{-\gamma}),$ with $P_{M}\coloneqq(\rho_{M}-\rho_{M_{b}})^{-1}$ as in (3.11). Then,
@@
where ^(^^$\boldsymbol{\hat{\prod}}^{\dagger}\coloneqq(\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}+\hskip-0.50003pt\gamma_{N}\hskip 0.50003pt\mathbb{I})^{\hskip-0.40974pt-1},$ $(M_{N})\subset\mathbb{N}$ and $(\gamma_{N})\subset(0,\infty)$ are sequences with $M_{N}\to\infty$ and $\gamma_{N}\rightarrow 0,$ and $\hat{h}_{1},\dots,\hat{h}_{\!M}$ are the eigenfunctions associated to the eigenvalues $\hat{\zeta}_{1}\geq\cdots\geq\hat{\zeta}_{M}\geq 0$ of $\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}.$ Further, $\hat{B}_{[i]},$ $B_{[i]},$ $\hat{\Psi}^{\prime\prime\prime}_{[i]},$ $\Psi^{\prime\prime\prime}_{[i]}\in\mathcal{S}_{\mathcal{H}^{\max(i,q)},\mathcal{H}}$ are for each $i,q$ defined by
@@
$\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\,\langle B_{q},h_{i}\otimes d_{j}\rangle^{2}_{\mathcal{S}}\,(1+i^{2\gamma}\!+\hskip-0.50003ptj^{2\gamma})<\infty,$
@@
There exists $d\in\mathbb{N}$ such that:
  • (a)
    The dimensions of all eigenspaces of $\boldsymbol{\prod}\boldsymbol{\prod}^{\ast}$ are bounded above by $d;$
  • (b)
    For all $N$ sufficiently large, with probability one the eigenvalues of $\boldsymbol{\hat{\prod}}\boldsymbol{\hat{\prod}}^{\ast}$ satisfy $\hat{\zeta}_{j}\neq\hat{\zeta}_{j+1}$ for all $j\in\{1,...,M+d\}.$
@@
[…]
@@ 6 Concluding remarks
where $(\varepsilon_{k})$ is i.i.d., $N$ is the sample size, and $k_{N}\to\infty$. This condition is weaker than standard invertibility (3.1), requiring only convergence in probability without assuming finite second moments or constraining $(k_{N})$. Standard invertibility implies mild Granger–Andersen invertibility, but not vice versa. As noted by Granger and Andersen (1978), certain functional time series, such as bilinear processes in separable Hilbert spaces (Hörmann and Kokoszka, 2010), satisfy mild but not standard invertibility. Others, like the non-linear MA(1)(1)( 1 ) process $X_{k}=\alpha_{1}(\varepsilon^{2}_{k-1})+\varepsilon_{k}$ in $L^{2}[0,1]$, may be stationary yet fail to be invertible in both senses.
@@
Appendix A Appendix